OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? 9x2 + 12x + 4

OpenStudy (whpalmer4):

Okay, are there any common factors in all three terms here?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Oh, what are they?

OpenStudy (anonymous):

2 & 3 ?

OpenStudy (whpalmer4):

Can you show me the result of factoring out those factors?

OpenStudy (anonymous):

3x + 2?

OpenStudy (whpalmer4):

well, no. We can't divide the terms by either of those numbers and get integer coefficients: 9/3 = 3 12/3=4 4/3= 4/3 9/2=9/2 12/2=6 4/2=2 There aren't any common factors to take out here. If the last term happened to be 3 instead of 4, or 6 instead of 4, then we would be able to factor out a 3, but that's not the case. \[9x^2+12x+4\]We'll use factoring by grouping. We multiply the coefficients of the first and third terms: \(9*4=36\) Now we factor that result, and look for a pair of factors which will add up to the coefficient of the middle term, \(12\). \[36*1\]\[18*2\]\[9*4\]\[6*6\]\[4*9\]\[2*18\]\[1*36\] \[6+6=12\], so we'll choose that. Now we rewrite your polynomial, splitting the middle term to use the two coefficients: \[9x^2+6x + 6x +4\]Now we group that into two groups:\[(9x^2+6x)+(6x+4)\]Now we factor each group:\[3x(3x+2)+2(3x+2)\]Notice that \((3x+2)\) is now a factor of each product term?

OpenStudy (whpalmer4):

We factor that again:\[3x(3x+2) + 2(3x+2) = (3x+2)(3x+2)\]

OpenStudy (anonymous):

3x + 1?

OpenStudy (whpalmer4):

No, \[(3x+2)(3x+2)\]is our factoring. You could also write that as\[(3x+2)^2\]

OpenStudy (whpalmer4):

Checking our work:\[(3x+2)(3x+2) = 3x*3x + 3x*2 + 2*3x + 2*2 =\]\[9x^2+6x+6x+4=9x^2+12x+4\checkmark\]